You are here: Foswiki>AdaptiveOptics Web>Logs>Log_20110610 (13 Jun 2011, JuanGuerra)Edit Attach

AO Commissioning with PISCES on 10 June 2011 UT

Observer(s): D. Miller, L Busoni, A. Puglisi, S. Esposito, R. Ragazzoni, K. Boutsia, J.C. Guerra (LBTO)

SW Support: (Tucson)

Telescope Support: J. Urban (LBTO)

Instrument Support: T. Shih(LBTO)

Summary

Details

3:30 Authorize irtc

4:25 iif_offset, 5, 0, COORD='DETXY' Moved star on Pisces up

4:27 iif_offset, 0, 5, COORD='DETXY' Moved star on Pisces left

4:41 iif_offset, 0, -8.0, COORD='DETXY' To put star near center of Pisces

4:42 iif_offset, 5.0, 0, COORD='RADEC' Moved star left on Pisces

4:42 iif_offset, 0, 50, COORD='RADEC' Moved star down on Pisces

4:44 iif_preset, M5_1396, MODE=ACE, GS=3, PA=-90., RS=0 Failed, AO Sup could not find Slope Null (wrong name, Alfio fixed)

4:45 iif_preset, M5_1396, MODE=ACE, GS=3, PA=-90., RS=0 Success

4:47 irc_runao Failed Too bright

5:00 AO guys work on closing the loop with Engineering GUI

5:15 Close the loop, but had a large coma.

5:30 Put Active Optics into Primary only, cleared active optics and C00, set secondary global offsets to values during the day (0, 1.7, 3.5, -313, -44, 0). I think this is a bad idea. The collimation lookup table from Andrew is good, and works properly with split mode.

6:30 Re-homed Pisces filter wheel and image was good. Out of focus. Juan Carlos and Konsantina had homed the FW during sky flats. This is the first time I have heard of a filter positioning error.

07:15 iif_preset, Pisces_5, MODE=ACQUIRE, GS=0, ROTMOD=IDLE Steve tweak IE, CA. ParAng = 90. Use equation PosAng = ParAng - 20

07:25 First out of focus imate pisces.20010608.1097.fits. Alfio ran the focus routine. Best focus is saturated even in FeII narrow band

08:13 iif_preset, Pisces_6, MODE=ACE, GS=8, RS=0, PA=90 Success

08:15 irc_runao

8:20 Sent a new presetTelescope and the AO Closed loop opened because of large tip (telescope moved) before a StopAO was received. Submitted IT#3434

8:30 iif_preset, BS9177, MODE=ACQUIRE, GS=0, ROTMOD=IDLE

8:33 iif_preset, BS9177, MODE=ACE, GS=13, RS=0, PA=180

8:35 irc_runao Success

9:00 AO guys done for a bit. JC and KB ready to collect J and K sensitivity data on BS9177

9:30 AO guys collect data on Pisces_7

10:00 iif_offset, 5.0, 0, COORD='DETXY' Failed with mirror panic. Not sure why. Sencond offset was successful, as were all the rest of the offsets.. We will try offsets with RADEC.


Pointing another magnitude star to close the loop to check loop performances, it has been seeing and strange behaves of the PSF in Pisces that it not expected, elongates stars, changing the focus didn't improve. The AO staff try to optimize the loop changing different Ao configurations with the same result of strange PSF. About 10:45 UT the three houses fun was turn off and the telescope was preset to a brighter magnitude start to close the loop, the result of the loop perfomances showed this strange shape of the PSF. After analysing different clope loop data, it was decided to home the filter wheel, and what happens now it the PSF changed in shape and it was enable to focus the star in Pisces and to get a close loop image with a strehl about 90%.???? (This number has to be check in a post-processing analysis)

-- DougMiller - 10 Jun 2011
Topic revision: r4 - 13 Jun 2011, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback